### v2

# HumanDreamer-X: Photorealistic Single-image Human Avatars Reconstruction via Gaussian Restoration

Abstract
Single-image human reconstruction is vital for digital human modeling applications but remains an extremely challenging task. Current approaches rely on generative models to synthesize multi-view images for subsequent 3D reconstruction and animation. However, directly generating multiple views from a single human image suffers from geometric inconsistencies, resulting in issues like fragmented or blurred limbs in the reconstructed models. To tackle these limitations, we introduce HumanDreamer-X, a novel framework that integrates multi-view human generation and reconstruction into a unified pipeline, which significantly enhances the geometric consistency and visual fidelity of the reconstructed 3D models. In this framework, 3D Gaussian Splatting serves as an explicit 3D representation to provide initial geometry and appearance priority. Building upon this foundation, HumanFixer is trained to restore 3DGS renderings, which guarantee photorealistic results. Furthermore, we delve into the inherent challenges associated with attention mechanisms in multi-view human generation, and propose an attention modulation strategy that effectively enhances geometric details identity consistency across multi-view. Experimental results demonstrate that our approach markedly improves generation and reconstruction PSNR quality metrics by 16.45% and 12.65%, respectively, achieving a PSNR of up to 25.62 dB, while also showing generalization capabilities on in-the-wild data and applicability to various human reconstruction backbone models.
Figure 1: Illustration of HumanDreamer-X. The pipeline initiates with a single-image reconstruction to generate a coarse 3D avatar, providing essential geometric and appearance priority for the restoration process. This approach facilitates the attainment of a higher-quality 3D avatar, suitable for subsequent animation tasks.
33footnotetext: Project Page: https://humandreamer-x.github.io/
1 Introduction
Creating 3D human avatars is gaining increasing significance across various domains, including virtual reality, gaming, and film production. Among the numerous methods, creating from a single image represents a common, practical, and user-friendly approach. Nevertheless, constructing a versatile human avatar with diverse shapes, appearances, and clothing from just a single image remains a substantial challenge.
Traditional reconstruction methods based on mesh, Neural Radiance Fields (NeRF) [nerf] and 3D Gaussian Splatting (3DGS) [3dgs] enable multi-view reconstruction but are incapable of achieving single-image reconstruction [SparseFusion, Text2Avatar, animatablegaussians, gaussianavatar, gauhuman, gart, instantavatar, humannerf].
Despite these advancements, single-image reconstruction remains particularly challenging. To address these limitations, current approaches for single-view human reconstruction often integrate techniques from image or video generation [sith, sifu, en3d, magicman, idol, charactergen, pshuman, anigs, gas]. A common strategy, as seen in methods like PSHuman [pshuman] and AniGS [anigs], involves first generating multi-view images using generative models [animateanyone, champ, mimicmotion, unianimate, animatex] and then performing subsequent reconstruction with mesh-based or Gaussian-based techniques. However, this decoupled paradigm heavily depends on the geometric and appearance consistency of the generative model. Any inconsistency in these aspects can lead to severe artifacts in the reconstruction stage, such as fragmented or distorted limbs.
To alleviate the aforementioned issues, we introduce HumanDreamer-X, a novel framework that integrates multi-view human generation and reconstruction into a unified pipeline. This integration facilitates mutual enhancement between the two processes, offering supplementary information on geometry and appearance. Consequently, this significantly improves the geometric consistency and visual fidelity of the reconstructed 3D models, effectively alleviating problems related to fragmented and blurred limbs in the generated models. As illustrated in Figure. 2, our framework first utilizes 3DGS to reconstruct the human body from a single image and then renders multi-view images. Leveraging the geometric representation capability of 3DGS, these rendered multi-view images provide strong geometric and appearance priority for the subsequent generative process. Building upon this, HumanFixer is introduced to refine the renderings, producing photorealistic images. Finally, these restored images are further utilized to guide 3DGS in reconstructing a high-quality human model. Compared to decoupled approaches, our unified framework effectively bridges the gap between reconstruction and generation, producing higher-quality avatars suitable for diverse downstream applications. Moreover, in the context of multi-view human generation, directly generating such videos often leads to blurriness because of temporal inconsistencies. To tackle this issue, we explore the inherent challenges associated with attention mechanisms within attention layers and propose a modulation strategy. This strategy effectively enhances geometric detail and identity consistency across multiple views, thus overcoming the limitations of traditional approaches.
Notably, our experiments demonstrate that our approach significantly improves the generation metrics by 16.45% in PSNR and the reconstruction metrics by 12.65% in PSNR. And also demonstrating its generalization capabilities on in-the-wild data and its adaptability to various human reconstruction backbone models.
The main contributions of this paper are summarized as follows:
  • We propose a novel framework, HumanDreamer-X, for generating animatable avatars from a single-view image by coupling 3D reconstruction with video restoration. This unified approach significantly enhances the quality and consistency of reconstructed avatars compared to existing decoupled methods.
  • We identify and address attention-related deficiencies in the generation of multi-view videos, which often lead to inconsistencies and blurriness. To mitigate these issues, we introduce an attention correction module that refines the temporal attention mechanism, thereby improving the quality of restored videos and ensuring better geometric and identity preservation.
  • Extensive experimental results demonstrate the superior performance of HumanDreamer-X across multiple datasets. Specifically, our method achieves higher fidelity in avatar reconstruction, showcasing its practical applicability and efficiency. It also demonstrates generalization capabilities on in-the-wild data and is applicable to various human reconstruction backbone models.
2 Related Work
2.1 Single-image Human Recosntruction
With the advancement of 3D representation methods such as mesh [Hu_2021_CVPR, meshanything], NeRF [nerf, mipnerf, zipnerf, mipnerf360, 10229247], and 3DGS [3dgs, mipgs, lu2024scaffold, deformablegs], human reconstruction has seen significant improvements [humanrf, hugs, 9787789, evahuman, avatarrex, 3dgsavatar, eva3d, humansplat, ag3d]. However, these methods typically require a large number of multi-view images or videos for reconstruction, which limits their applicability. In contrast, Single-image human reconstruction is a more flexible task, aiming to reconstruct a 3D human model from just one image [anigs, pshuman, charactergen, idol, sith, sifu, magicman, en3d]. However, it is inherently an ill-posed problem, presenting considerable challenges.
Current single-image human reconstruction methods integrate techniques from generative approaches [sd, svd, controlnet, yang2024cogvideox, zhu2024sora]. These methods first generate unseen viewpoints [sith] or multi-view images [anigs, pshuman, en3d, magicman, charactergen], followed by employing 3D representations to reconstruct the human. To ensure geometric consistency, methods such as SMPL [anigs, pshuman], masks [sith], segmentation maps [magicman], and pose estimation [charactergen] are utilized to drive the generation process. However, these generate-then-reconstruct approaches heavily rely on the geometric and appearance consistency of the generated images, and inconsistencies can lead to issues like fragmented limbs and blurriness in the reconstructed models. Recently, there have been attempts [idol] to infer 3DGS parameters directly from a single image using feed-forward networks, but this requires extensive datasets for training. The approach most closely related to ours is SIFU [sifu], which refines the coarse texture of reconstructed models using generative models. However, SIFU uses a frozen image generation model for frame-by-frame refinement, making it difficult to maintain inter-frame consistency.
Figure 2: Overall framework of the proposed HumanDreamer-X. The process begins by initializing a coarse 3DGS avatar using a reference image. A rendered video serves as a , providing geometric and appearance priors. Subsequently, HumanFixer performs video restoration, wherein an attention modulation is employed to enhance video consistency. Throughout this process, the restored video is used to continuously update the 3DGS model, ultimately resulting in a refined 3DGS avatar.
2.2 Controllable Human Generation
With the rapid advancements in image and video generation [svd, controlnet, yang2024cogvideox, ma2024latte, eg3d, worlddreamer, yan2021videogpt, videoldm], diffusion-based human generation models [ipadapter, animateanyone, champ, mimo, animateanyone2, mimicmotion, animatex, emo, followyouremoji, aniportrait, 10543061] have gained increasing attention. Among these, Animate Anyone [animateanyone] stands out as a representative work, which constructs a 3D U-Net upon Stable Diffusion [sd] for modeling temporal information, and incorporates skeleton information as driving signals for controlling motion generation.
To further enhance temporal consistency, MimicMotion [mimicmotion] leverages pre-trained video generation model SVD [svd] to improve the coherence of generated sequences. In contrast, UniAnimate [unianimate] employs MAMBA-based [mamba] techniques for enhanced temporal modeling. Additionally, several approaches explore different motion control signals. For instance, DisCo [disco] and MagicAnimate [magicanimate] utilize DensePose [densepose] as a representation of the human body [disco, magicanimate], while Champ [champ] integrates multi-modal information including depth, normal, and semantic signals derived from the 3D parametric human model SMPL [smpl]. Despite their ability to generate photorealistic frames, these models often struggle to fully ensure geometric and appearance consistency across frames. This inconsistency poses significant challenges for subsequent reconstruction tasks.
3 Method
3.1 Preliminary
3D Gaussian Splatting. 3DGS represents a voxel-based rendering technique for scene representation, where the core concept involves modeling the scene using an optimized set of 3D Gaussian distributions $\mathcal{G}=\{\mathcal{N}(\mathbf{x}_{i},\Sigma_{i})\}_{i=0}^{N}$. Each Gaussian distribution $\mathcal{N}$ is parameterized by its position $\mathbf{x}_{i}\in\mathbb{R}^{3}$, covariance matrix $\Sigma_{i}$ (which defines the ellipsoidal shape), and radiance attributes such as color $\mathbf{c}_{i}$ and opacity $\alpha_{i}$. During rendering, these Gaussian distributions are projected onto the image plane. The contribution of each pixel $\mathbf{u}$ is computed via radial basis functions defined as $w_{i}(\mathbf{u})=\alpha_{i}\exp\left(-\frac{1}{2}(\mathbf{u}-\mathbf{u}_{i})^{\top}\Sigma_{i}^{-1}(\mathbf{u}-\mathbf{u}_{i})\right)$, where the weight $w_{i}$ decays with distance from the center of the Gaussian. This process is achieved through differentiable rasterization, which integrates depth sorting and weighted blending ($\sum_{i}w_{i}\mathbf{c}_{i}/\sum_{i}w_{i}$), thereby facilitating gradient-based optimization for high-quality scene reconstruction.
Video Diffusion Model. Video Diffusion Models (VDM) are generative frameworks that extend diffusion-based image synthesis to dynamic scenes by modeling sequential data in a latent space. The core principle involves a forward diffusion process that gradually corrupts video data $\mathbf{x}_{0}$ into noise via $T$ steps: $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$, where $\beta_{t}$ controls noise injection. VDM enhances temporal coherence by integrating spatiotemporal layers into a pre-trained latent space, enabling high-fidelity video generation from images or text while leveraging efficient VAE-based compression. As a reverse denoising network it learns to approximate the posterior $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, typically parameterized as $\mathbf{x}_{t-1}=\mu_{\theta}(\mathbf{x}_{t},t)+\sigma_{t}\epsilon$ ($\epsilon\sim\mathcal{N}(0,\mathbf{I})$), reconstructing coherent video frames through iterative refinement.
3.2 Overall Framework of HumanDreamer-X
Traditional human reconstruction methodologies [animatablegaussians, gaussianavatar, gauhuman, gart, instantavatar, humannerf] encounter substantial challenges due to their reliance on multi-view images. Recent advances [sith, sifu, en3d, magicman, idol, charactergen, pshuman, anigs, gas] have endeavored to address this limitation by leveraging generative priors to compensate for the absence of invisible views. Nevertheless, the decoupling of generation from reconstruction has resulted in a deficiency of geometric consistency among the generated multi-view images. Our proposed framework, HumanDreamer-X, integrates reconstruction and generation into a unified pipeline. In this pipeline, the reconstruction step provides initial geometry and appearance priorities, and the generative model then refines the reconstructions by restoring coarse renderings. The overall framework is illustrated in Fig. 2.
Specifically, we first use the 3DGS model [animatablegaussians, gaussianavatar] to reconstruct an avatar $\mathcal{A}_{c}$ from a single reference image $\mathcal{I}_{R}$:
$\mathcal{A}_{c}=\text{3DGS}(\mathcal{I}_{R},\theta_{\text{SMPL}}),$ (1)
Following previous avatar reconstruction models, the initial point cloud for 3DGS is derived from the estimated Skinned Multi-Person Linear (SMPL) $\theta_{\text{SMPL}}$. This ensures that even a single image can reconstruct coarse human geometry and appearance, providing a basic priority for subsequent refinement. Then, a multi-view video $\mathcal{V}_{c}$ of the avatar is rendered using the trained 3DGS avatar:
$\mathcal{V}_{c}=\{\mathcal{A}_{c}(d_{i})\mid i=1,2,\ldots,n\},$ (2)
where $d_{i}$ represents a specific horizontal angle, and $\mathcal{A}_{c}(d_{i})$ renders the avatar at that angle. Due to the limited availability of only one viewpoint, the resulting model $\mathcal{A}_{c}$ provides only basic geometric and appearance priority of the avatar, leaving issues such as blurriness and artifacts in unseen views unresolved. To address these issues, we introduce HumanFixer, a video generation model designed to restore details in the initial 3DGS renderings.
HumanFixer is built upon the pretrained video diffusion model [svd], utilizing the coarse video $\mathcal{V}_{c}$ and the reference image $\mathcal{I}_{R}$ as conditions to generate multi-view refined videos (see Sec. 3.3 for more details). The refined videos $\mathcal{V}_{r}$ capture the textures and geometry present in the reference image $\mathcal{I}_{R}$, making them suitable for modeling a refined human avatar $\mathcal{A}_{r}$.
This approach leverages the geometry and appearance priority provided by 3DGS and exploits the temporal consistency inherent in video generation models, ensuring enhanced multi-view consistency in the repaired avatar. Additionally, to address blurriness in multi-view video generation, we investigate attention mechanisms and propose an attention modulation strategy (see Sec. 3.4 for more details). The refined video, enhanced through this strategy, is then utilized to optimize the human avatar.
3.3 Training and Inference of HumanFixer
Reconstructed avatars derived directly from single-image inputs exhibit issues such as blurring at invisible views. To address these problems, we introduce HumanFixer for refining coarse avatars. This section will present the methodologies for the training and inference of HumanFixer.
Figure 3: The creation of the dataset for training HumanFixer. First, we use Blender to render scans and obtain the ground truth video. Next, we employ the frontal image and its corresponding SMPL prior to reconstruct a coarse 3DGS model, followed by rendering multi-view videos. This process yields paired video data for training.
Training. HumanFixer hinges on constructing a coarse-refined pair dataset. In this section, we propose a concrete pipeline for dataset collection, as illustrated in Fig. 3. The steps are as follows: Initially, we utilize Blender to render a multi-view video $\mathcal{V}_{\text{gt}}$ from each 3D scan, serving as the ground truth video. Subsequently, we leverage the GS model to reconstruct human avatar from single image, and use Eq. 2 to render multi-view coarse images $\mathcal{V}_{c}$, paired with their corresponding ground truth videos $\mathcal{V}_{\text{gt}}$, form the repair dataset. This dataset is designed to refine avatar reconstructions by providing examples of both low-quality and high-quality renderings, allowing the HumanFixer model to learn how to enhance the coarse images into refined, detailed videos.
The architecture of HumanFixer is illustrated in Fig. 2. It employs SVD [svd] as the backbone and leverages the low-quality video to guide the generation process. Additionally, it integrates a CLIP [clip] encoder to inject reference image information through cross-attention mechanisms.
During the training of HumanFixer, we first feed the coarse video $\mathcal{V}_{c}$ into a Variational Autoencoder (VAE) [vae] encoder $\mathcal{E}$ to obtain the latent feature $z_{c}=\mathcal{E}(\mathcal{V}_{c})$. Then, $z_{c}$ is used as a condition, providing both geometric and appearance priority. It is concatenated with $z_{gt}=\mathcal{E}(\mathcal{V}_{gt})$, serving as the input to the model. To maintain identity consistency across multiple views, we utilize a reference image $\mathcal{I}_{R}$ as a source of identity information. The face embedding of the reference image $\mathcal{M}_{f}\in R^{n_{t}\times C}$ is extracted using an existing facial embedding extraction model from $\mathcal{I}_{R}$. $n_{t}$ denotes number of tokens and $C$ means dimension of cross-attention. The model’s output is:
$\epsilon_{\text{target}}=h_{\theta}(z_{gt}^{t},z_{c},\mathcal{M}_{f}),$ (3)
Where $h_{\theta}$ denotes the HumanFixer module with parameters $\theta$, $z_{gt}^{t}$ represents the latents of the ground truth video $\mathcal{V}_{gt}$ at the noising time step $t$, and $z_{c}$ signifies the latents of the coarse video $\mathcal{V}_{c}$. As described in [svd], we use a predicted target loss [svd] for optimizing the HumanFixer model.
Inference. After training HumanFixer, the coarse video $\mathcal{V}_{c}$ to be refined and the reference image $\mathcal{I}_{R}$ are used as inputs to obtain the refined video $\mathcal{V}_{r}$.
$\mathcal{V}_{r}=\text{EDM}(h_{\theta}(z^{T},z_{c},\mathcal{M}_{f})),$ (4)
where $z^{T}$ is the initial latent noise, and we use EDM scheduler [svd] for denoising.
Figure 4: Attention weights visualization. The left and right sides show the head 0 attention weights at the temporal self-attention stage for training on cyclic videos without and with an attention modulation, respectively. Brighter colors indicate higher weights.
Table 1: Multi-view video generation comparison of other SOTA methods. Bold indicate the best result.
Method Testset PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
PSHuman [pshuman] CustomHumans 21.998 0.826 0.1945 103.808
Champ [champ] 20.228 0.889 0.2438 115.031
HumanFixer (Ours) 25.618 0.882 0.0687 87.149
Champ [champ] THuman2.1 17.547 0.859 0.2701 129.629
HumanFixer (Ours) 23.741 0.889 0.0720 94.570
3.4 Attention Mechanism Analysis
Generating multi-view videos differs from generating standard videos due to the distinct temporal relationships involved. In typical videos, adjacent frames are strongly correlated, with correlation decreasing as the distance between frames increases. However, in multi-view videos, which encompass a full circle of views, the final frames have a strong relationship with the initial frames. Given that our model is fine-tuned on SVD, which is pretrained on standard videos, the pretrained parameters align more closely with the assumption of strong correlations between adjacent frames.
For a multi-view video with a total of $N$ views, we define a non-cyclic video as one where each frame corresponds sequentially to views 0,1,,N10,1,...,N-1. Conversely, a cyclic video involves frames corresponding to views 0,1,,N1,00,1,...,N-1,0, effectively looping back to the initial view.
Compared to non-cyclic videos, cyclic videos append the first frame to the end, making the training data more suitable for models pretrained on the assumption of strong correlations between adjacent frames. This should theoretically enhance overall consistency. However, during training, we observed that directly training on non-cyclic videos allows view 0 to develop stronger associations with views N1N-1, N2N-2, and N3N-3. Nevertheless, a significant discontinuity was noted between view N1N-1 and view 0, whereas the difference between view $1$ and view 0 was relatively minor.
To address this issue, we further analyzed the temporal self-attention mechanism within the model. As illustrated in the left panel of Fig. 4, the attention weights for the starting and ending frames are significantly higher than those for intermediate frames. We hypothesize that this occurs because, in cyclic videos, the first and last frames are identical, leading to naturally higher attention weights compared to those calculated between different frames. This phenomenon suppresses information flow among intermediate frames.
In summary, while cyclic videos aim to improve consistency by leveraging adjacency assumptions, they introduce challenges related to discontinuities and uneven attention weight distribution, necessitating further refinement of the temporal self-attention mechanism. To mitigate this issue, we apply an attention modulation, as shown in Fig. 2, which ensures that the first and last frames do not receive attention from the model, either from themselves or from each other.
Formally, let the attention mask $M\in\mathbb{R}^{(N+1)\times(N+1)}$ be defined as:
$M(i,j)=\begin{cases}-\infty,&\text{if }(i,j)=(0,0),(0,N),(N,0),(N,N)\\ 0,&\text{otherwise}.\end{cases}$
Then, this attention mask is added to every temporal self-attention module in the model:
$\text{Attn}(Q,K,V,\mathcal{M})=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}+\mathcal{M}\right)V,$ (6)
where $Q$ for query, $K$ for key, $V$ for value, $d_{k}$ the dimension for scaling down the dot product results, $\mathcal{M}$ denotes the attention mask on temporal self attention.
This formulation effectively suppresses attention weights for the first and last frames, preventing the model from being unduly influenced by these boundary frames during video generation. Fig. 4 illustrates that, after training on cyclic videos with the attention modulation, the attention weights shift from being predominantly focused between the 0th and $N$th frames to resulting in a more evenly distributed attention pattern.
4 Experiments
This section outlines our experimental framework, including the datasets, implementation specifics, and evaluation criteria. We then provide both quantitative and qualitative results to highlight the outstanding performance of the proposed HumanDreamer-X.
Figure 5: Comparison of generation with SOTA methods. Note that PSHuman’s training dataset contains all of the CustomHumans. Best viewed with zoom-in.
4.1 Experiment Setup
Dataset. Our experiments are conducted on a variety of datasets. To ensure that the human subject occupies a significant portion of the frame during training, we filter out instances where the arms are excessively spread, which would otherwise result in an overabundance of background content. The final training set comprises 388 scans from CustomHumans [customhumans] and 1929 scans from THuman2.1 [function4d]. For testing, we use 39 samples from [customhumans], 32 samples from [function4d]. All training data is standardized to a resolution of $960\times 640$, with cyclic video sequences consisting of 19 frames (18 multi-view frames plus the first frame repeated).
To ensure the reconstruction of a complete avatar, it is stipulated that the input to HumanFixer must include facial information as identity cues. For the THuman2.1 dataset [function4d], facial detection is performed using InsightFace [arcface], and the image with the largest facial area is designated as the reference image. For the CustomHumans dataset [customhumans], the first frame is directly selected as the reference image.
Baselines. For multi-view video generation, we utilize PSHuman [pshuman] and Champ [champ] as baselines. For 3D avatar reconstruction, we compare our method with a variety of existing single-image human reconstruction approaches. These include mesh-based methods such as PIFU [pifu], PaMIR [pamir], SiTH [sith], and PSHuman [pshuman], as well as recent feed-forward 3D Gaussian Splatting methods like IDOL [idol] and LHM [lhm]. In contrast to these end-to-end approaches, our framework adopts a two-stage generation-then-reconstruction pipeline. To validate the effectiveness of our framework across different generation strategies, we use either Champ [champ] or HumanDreamer-X for the initial human prior generation stage, while keeping the subsequent 3D reconstruction stage fixed—implemented on two distinct 3DGS backbones: Animatable Gaussians [animatablegaussians] and GaussianAvatar [gaussianavatar].
4.2 Main Results
Quantity comparison results on multi-view generation. We compare the multi-view generation performance of HumanFixer with several SOTA methods, and the results are presented in Tab. 1. The findings indicate that HumanFixer achieves superior video consistency. Fig. 5 visually illustrates the enhanced multi-view consistency produced by our method.
Table 2: 3D reconstruction comparison. * denotes the metric is from PSHuman[pshuman].
Method Testset PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
PSHuman CustomHumans 20.089 0.8439 0.1770 87.816
LHM 20.153 0.9122 0.1121 110.341
IDOL 18.333 0.9209 0.0882 187.710
Champ with GaussianAvatar 19.673 0.8789 0.2643 164.554
HumanDreamer-X with GaussianAvatar 23.639 0.9100 0.2427 114.804
Champ with Animatable gaussians 16.853 0.9157 0.1251 122.752
HumanDreamer-X with Animatable gaussians 22.631 0.9458 0.0729 71.250
PIFU THuman2.1 19.3957 0.8327 0.1001 -
PaMIR 19.4130 0.8324 0.0988 -
SiTH 18.458 0.8200 0.1004 -
PSHuman 20.855 0.8636 0.0764 -
LHM 19.275 0.8913 0.1062 139.566
IDOL 19.348 0.9321 0.0919 201.377
Champ with GaussianAvatar 18.264 0.8842 0.2639 129.413
HumanDreamer-X with GaussianAvatar 19.328 0.8945 0.2578 132.200
Champ with Animatable gaussians 18.908 0.9328 0.1278 176.836
HumanDreamer-X with Animatable gaussians 21.091 0.9403 0.0968 78.174
Figure 6: Comparison of 3D reconstruction with SOTA methods. Best viewed with zoom-in.
Quantity comparison results on 3D reconstruction. To evaluate the 3D reconstruction quality, we render the reconstructed models from 18 different views and calculate the PSNR, LPIPS, SSIM, and FID metrics. We compare our approach against various baselines to demonstrate the broad applicability of our proposed framework. As shown in Tab. 2, our framework supports the integration of more advanced 3DGS backbones for enhanced reconstruction performance, thereby achieving superior results. We also conducted a visualization comparison experiment, as shown in Fig. 6. Compared to other methods, ours delivers higher detail fidelity while mitigating issues like poor identity consistency, missing hands, and limb sticking artifacts. Specifically, compared to the second-best method, PSHuman, our approach shows improvements in PSNR, SSIM, LPIPS and FID by 12.65%, 12.07%, 58.81% and 18.86% on CustomHumans.
Table 3: Ablation Study on the Use of Coarse 3DGS.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
w/o coarse 3DGS 16.824 0.621 0.3122 170.659
w/ coarse 3DGS 25.618 0.882 0.0687 87.149
Table 4: Ablation Study on Attention Modulation in the Generation Stage on the CustomHumans Subset.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
w/o attention mask (non-cyclic) 25.514 0.885 0.0704 95.065
w/o attention mask (cyclic) 25.667 0.800 0.1453 106.705
w attention mask (cyclic) 25.618 0.882 0.0687 87.149
Table 5: Ablation Study on Attention Modulation in the Reconstruction Stage using Animatable Gaussians [animatablegaussians].
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
w/o attention mask (non-cyclic) 21.309 0.9398 0.0867 112.812
w/o attention mask (cyclic) 20.867 0.9351 0.0955 139.693
w attention mask (cyclic) 22.631 0.9458 0.0729 71.250
4.3 Ablation Study
We conduct an ablation study to evaluate the importance of coarse 3DGS-based multi-view reconstruction as a conditioning signal during training. Specifically, we train the generation model without using the coarse 3DGS rendering as input condition. As shown in Tab. 3, the significant performance drop in all metrics, demonstrates the critical role of this geometric prior in guiding high-fidelity human video generation. These results validate that coarse 3DGS rendering provides effective structural supervision, significantly improving the quality and consistency of the generated multi-view videos.
Furthermore, we train HumanFixer on the CustomHumans dataset [customhumans], comparing non-cyclic videos (18 frames) and cyclic videos (19 frames) with and without attention modulation, under both multi-view video generation (see Tab. 4) and 3D reconstruction settings (see Tab. 5). Experimental results show that while video quality degrades when transitioning from non-cyclic to cyclic inputs, the incorporation of attention modulation notably improves performance, achieving the best overall results. This confirms the effectiveness of our proposed attention mechanism in enhancing the fidelity and temporal coherence of generated avatar sequences.
5 Limitations and Conclusions
Limitations. Our method follows a two-stage pipeline: it first reconstructs a coarse human avatar, followed by restoration and final refinement. As described in the paper, this approach results in relatively high computational cost. Depending on the chosen 3DGS baseline, each full human avatar modeling process takes approximately 7 to 15 minutes, with the reconstruction stage alone accounting for about 90% of the total time. Employing faster reconstruction techniques could help alleviate this computational bottleneck.
Conclusions. Single-image human reconstruction is vital for digital human modeling but remains challenging due to geometric inconsistencies and limited visual fidelity. To address these issues, we propose HumanDreamer-X, a unified framework for multi-view human generation and reconstruction. It leverages 3DGS for robust initialization and introduces HumanFixer to refine geometry and appearance, along with an attention modulation strategy to enhance multi-view consistency. Experimental results show that our method significantly improves reconstruction quality, generalizes well to in-the-wild data, and is compatible with various backbone models.